OpenStudy (anonymous):

Simplify, show work please √18 ---- √54

OpenStudy (anonymous):

1/Sqrt(3)

OpenStudy (radar):

\[3\sqrt{2}/3\sqrt{6}\]

OpenStudy (radar):

\[3\sqrt{2}/3\sqrt{3}\sqrt{2}\]

OpenStudy (radar):

cancel out common factors that exist in both numerator and denominator. Like the 3, and the sqrt 2 This leaves 1/\[1/\sqrt{3}\] This is not considered "good form" (a radical in the denominator so multiply both numerator and denominator by sq rt of 3 \[\sqrt{3}/3\] or 1/3 sqrt 3

OpenStudy (radar):

did you follow, I tried to show work.

OpenStudy (radar):

I am standing by for any questions or clarification

OpenStudy (radar):

I just noted that I did not show how the original numbers were changed. 18 is the same as 9 x 2. Well 9 is a perfect square with a square root of 3 which can be removed from the radical leaving only the 2 in the radical. Comphrendo?

OpenStudy (radar):

the 54 is the same as 9 times 6 and the same procedure can be followed leaving only the 6 inside the radical.